OpenStudy (unklerhaukus):

\[{dy \over dx}={ x-y+2 \over x-y+1}\]

OpenStudy (unklerhaukus):

\[{dy \over dx} ={( X+q)-(Y+p)+2\over ( X+q)-(Y+p)+1}\] \[{dy \over dx} ={ X-Y +(q-p+2)\over X-Y+(q-p+1)}\]

OpenStudy (unklerhaukus):

but i cant sole these simultaneously \[q-p+2=0\]\[q-p+1=0\]there are no solutions /....?